# Unified Theory of Magnetization Temperature Dependence in Ferrimagnetics

### v1

@@ Abstract
Recent advancements in spintronics and fundamental physical research have brought increased attention to the rare-earth-based magnetically ordered materials. One of the important properties of these materials is the temperature dependence of the spontaneous magnetization $M(T)$. Recently, a successful framework was proposed for the theoretical description of $M(T)$ across the entire temperature range from $T=0$ K to the Curie temperature $T_{{}_{\rm C}}$ in simple cubic ferromagnetics, EuO and EuS. We extend this approach to compute and analyze $M(T)$ for a more complex magnetic material: multi-sublattice collinear ferrimagnetics such as Yttrium Iron Garnet, (YIG) Y3Fe5O12. YIG is a unique material that has a high Curie temperature well above room temperature. It features a cubic crystallographic structure characterized by low anisotropy and minimal spin-orbit interaction. These properties result in an exceptionally low spin-wave damping, which is important for fundamental studies as well as engineering applications. The primary challenge in describing ferrimagnetics theoretically arises from its intricate magnetic structure, characterized by multiple magnetic sublattices. We analyzed and generalized for multi-sublattice collinear ferrimagnetics two well-known approximations describing $M(T)$. The first approach is the Bloch $\frac{3}{2}$-law, which describes the suppression of $M(T)$ due to spin-wave excitation, and is valid in the low-temperature limit $T\ll T_{{}_{\rm C}}$. The second one is Weiss’s mean-field approximation, which provides a reasonable description of $M(T)$ near $T_{{}_{\rm C}}$. Using a single tuning parameter, we combine these two approaches to describe $M(T)$ for any $0\leq T\leq T_{{}_{\rm C}}$. The theoretical result for $M(T)$ aligns well with our measurements and the previously available experimental data across the entire temperature range. We also demonstrate that experimental and theoretical dependences $M(T)$ follow the mean-field prediction $\sqrt{T_{{}_{\rm C}}-T}$ for almost all temperatures.
Figure 1: (a) The crystallographic structure of YIG features a body-centered cubic (BCC) unit cell with the Oh symmetry group, occupying half of a cube. This structure includes four formula units (Y3Fe5O12)×$\times 4$, amounting to a total of 80 atoms. Among these, there are 20 magnetic iron ions (Fe3+) divided into two groups: 8 Fe3+ ions occupying the (a) octahedral sites, represented as a blue ball and a blue-shaded area, and the remaining 12 Fe3+ ions located in (d) tetrahedral sites, represented as a green ball and yellow-shaded area. The dodecahedra (c) site occupied by Yttrium ions is shown as a black ball and green-shaded area, the O2- ions are drawn in red.
@@ I Introduction
Promising perspectives for engineering applications revived the attention to fundamental studies of the physics of magnetic materials. To our understanding, the first step in developing a consistent theory of magnetic dielectrics is the study of their most fundamental properties: the temperature dependence of spontaneous magnetization $M(T)$ in thermodynamic equilibrium. Recently [6], the behavior of $M(T)$ was studied for the simplest possible magnetic system: cubic ferromagnetic crystals containing only one magnetic ion per elementary cell. Currently, the only known and experimentally studied example of such a system is body-centered crystals of EuO and EuS.
The analysis in [6] is based on the well-known Weiss mean-field approximation (MFA) [7] with the Heisenberg exchange interaction [8]. In the MFA, an effective magnetic field acts on each magnetic ion due to exchange interactions with its nearest neighbors. The larger their coordination number $Z$ (i.e., the number of the nearest neighbours), the smaller the fluctuations of this field, and the better the MFA performs. In the limit $Z\to\infty$, MFA becomes exact. In EuO, $Z=12\gg 1$ and the resulting value of the Curie temperature for EuO $T^{{}^{\rm MF}}_{{}_{\rm C}}\approx 86.6\,$K, predicted by MFA, is only about 20% larger than its experimental value $T^{\rm Exp}_{{}_{\rm C}}\approx 69.8\,$K.
@@
A natural further step in the development of the theory of the spontaneous magnetization is to extend it to multi-sublattice ferri- and antiferromagnetics. This is the objective of the present paper. To make our study more concrete and provide a basis for comparing our theory with experiments, we have chosen a specific example: the twenty-sublattice ferrimagnetic Yttrium Iron Garnet (YIG), Y3Fe5O12. YIG is one of the most important materials in microwave magnetics. In physics of magnetically ordered crystals, it plays a crucial role much like the role of Drosophila melanogaster (fruit fly) in genetics [9] or water in hydrodynamics  [10].
@@
Various experimental datasets and theoretical results are labelled as follows. Experimental data: blue dots ① – 1964-Anderson results [20], orange dots ② – 1974 data Hansen and others [21], black solid line ③ – our current results of a 1 mm YIG sphere magnetized in the ⟨111⟩ direction were measured via VSM.
Violet dashed line ④ – all-temperature fit of experimental data ② $\overline{S}^{z}=A\cdot(T_{{}_{\rm C}}-T)^{\beta}$ with $A$ = 1.045 and $\beta\simeq 0.50\pm 0.005$; for details see Appendix F.
@@
Our aim in this paper is to improve the description of multi-sublattice ferrimagnetics, focusing on a theory of spontaneous magnetization $M(T)$ over an entire temperature range from $T=0$ to the Curie temperature $T_{{}_{\rm C}}$.
@@
We begin our study by presenting in Fig. 2 the earlier available experimental data [20, 21] for $\overline{S}^{z}(T)/\overline{S}^{z}(0)$, as blue and orange dots labeled as ① and ② and our own accurate measurements down to temperatures near 2 K, marked as the black solid line ③. The experimental Curie temperature $T_{{}_{\rm C}}^{\rm Exp}\approx 560\,$K is clearly visible in the figure. A small tail above $T_{{}_{\rm C}}^{\rm Exp}$ originates from the external magnetic field (about 1 T) used in experiments to suppress the domain structure of YIG, interfering with measurements.
@@
[…]
@@
To this end, in Sec. III.1 we consider the MFA for two-sublattice ferrimagnets and derive Eqs. (15) that describe it. Next, we derived an analytical expression for the Curie temperature, Eq. (16a), along with the ratio of the sublattice magnetizations near $T_{{}_{\rm C}}$, Eq. (16b). The numerical solution of Eq. (15) in the form Eq. (8) is plotted in Fig. 2 by the blue solid line ⑤. As expected, near $T_{{}_{\rm C}}^{{}^{\rm MF}}$ it demonstrates normal scaling behavior $\overline{S}^{z}(T)\propto\sqrt{T_{{}_{\rm C}}^{{}^{\rm MF}}-T}$, denoted by the blue dashed line ⑥.
The resulting Curie temperature $T_{{}_{\rm C}}^{{}^{\rm MF}}\approx 894\,$K calculated using values of the YIG exchange integrals [23] exceeds its experimental value $T_{{}_{\rm C}}^{\rm Exp}=560\,$K almost twice, while in EuO the discrepancy is only about 20%. This is discouraging but not unexpected, since the dominant inter-sublattice interactions in YIG are carried through only $Z_{\rm YIG}\approx 5$ neighbors on average. Therefore, the applicability parameter $51/Z_{\rm YIG}\approx 1/$ 5 is larger than $121/Z_{\rm EuO}\lesssim 1/$ 12, and poorer performance of MFA compared to EuO is expected.
@@
This issue is addressed in Section III.2 where we formulated spin-wave approximation (SWA) Eq. (21) for suppression of magnetization in multi-sublattice magnetics. To calculate the magnitization in the framework of SWA, an explicit knowledge of the magnon frequency spectra is required.
[…]
@@
Formally speaking, Eq. (21) is exact in the low-temperature range, say for $T\ll T_{{}_{\rm C}}$. To our surprise, the numerical solution of Eq. (21), represented by the green solid line ⑦ in Fig. 2, aligns well with the experimental data ①, ②, and ③ from $T=0$ up to $T\approx 400\,K$. At higher temperatures $\overline{S}^{z}(T)$ descreses almost linearly, crossing zero at $T_{\times}\approx 770\,$K and then formally becomes negative.
[…]
@@
In the simple case of a one-sublattice ferromagnet, this concept was demonstrated in Ref. [6] and briefly reminded in Sec. IV.1 for completeness. Its generalization for multi-sublattice ferrimagnetics like YIG is not straightforward and requires significant efforts, including the application of the diagrammatic perturbation approach for spin operators, such as the Belinicher-Lvov diagrammatic technique [24]. This approach is beyond the scope of the current paper.
@@
The numerical solution of Eqs. (27) and (30) with $\delta=0.428$ is shown in Fig. 2 as the red solid line ⑧. Bearing in mind that we computed the solution ⑧ using a single tuning parameter $\delta$, we consider its agreement with experimental data for $\overline{S}^{z}(T)$ across the entire temperature range as fully satisfactory.
It is worth noting that the theoretical dependence ⑧ of $\overline{S}^{z}$ retains normal scaling over $100\lesssim T\lesssim T_{\rm{}_{C}}$, much wider than the MFA.
We summarize our findings in Section V.
@@ II.1 Weiss-Heisenberg mean-field approximation
The analysis of magnetic properties in magnetodielectrics begins with the Weiss mean-field approximation [7] (MFA) with the Heisenberg exchange Hamiltonian [8]. In our paper, we adopt the simplest form of this Hamiltonian, which describes the interaction between the spins $\bm{S}_{i}$ and $\bm{S}_{j}$, located at positions $\bm{r}_{i}$ and $\bm{r}_{j}$, respectively:
@@
Here, exchange integrals $J_{ij}$ represent the energy associated with the exchange interactions.
@@
$~{}\hskip-14.22636ptb_{{}_{S}}(x)$ $=$ $\Big{(}S+\frac{1}{2}\Big{)}\coth\!\Big{[}\Big{(}S+\frac{1}{2}\Big{)}x\Big{]}-\frac{1}{2}\coth\Big{(}\frac{x}{2}\Big{)}\ .$ (4b)
@@ II.2 Crystallographic and magnetic structure of YIG
The first obstacle in our pursuit is YIG’s complicated crystallographic structure. The unit cell of YIG includes four formula units (Y3Fe5O12)×$\times 4$, i.e., 80 atoms. The lattice has a body-centered cubic (BCC) unit cell of Oh symmetry group with a lattice constant of $a=12.4\times 10^{-8}$ cm, occupying half of a cube as shown in Fig. 1. Magnetic Fe3+ ions have spin $S=\dfrac{5}{2}$ and magnetic moment $\mu=5\,\mu_{{}_{\rm B}}$, where $\mu_{{}_{\rm B}}$ is the Borh magneton. They occupy two inequivalent positions with regards to the symmetry of their immediate O2- environment - $N_{\rm a}=8$ octahedral sites (a), shown in Fig. 1 as a blue ball and a light-blue-shaded area, and $N_{\rm d}=12$ tetrahedral sites (d), shown as a green ball and a yellow-shaded area. Explicit coordinates of magnetic Fe3+ ions are listed in Appendix B.
@@
Over several decades, a variety of exchange integrals for YIG have been proposed, some of which are presented in Tab. 1 and discussed in Appendix C. For this paper, we adopt the following values:
@@ II.3 Spin-waves spectra of YIG
$\begin{split}H(\bm{k})=&\sum_{i,j=n_{+}+1}^{n}A_{ij}(\bm{k})a_{i}^{*}(\bm{k})a_{j}(\bm{k})\\ &+\sum_{i,j=1}^{n_{+}}D_{ij}(\bm{k})a_{i}^{*}(\bm{k})a_{j}(\bm{k})\\ &+\sum_{i=n_{+}+1}^{n}\sum_{j=1}^{n_{+}}\big{[}B_{ij}(\bm{k})a_{i}^{*}(\bm{k})a_{j}^{*}(-\bm{k})+\mbox{c.c}\big{]}\,,\end{split}$ (10)
Recall that in YIG $n_{+}=12$, $n_{-}=8$, and $n=n_{+}+n_{-}=20$. The explicit and quite combersome equations for $A_{ij}(\bm{k})$ $B_{ij}(\bm{k})$ and $D_{ij}(\bm{k})$ one finds in [23].
@@ III.1 MFA for two-sublattice ferrimagnets
As discussed in the Introduction, the MFA focuses solely on the mean values of the spin $\overline{S}^{z}_{j}$ at site $\bm{r}_{j}$. In YIG, there are only two crystallographically distinct positions of Fe ions, labeled as (a) and (d), with mean spins $\overline{S}^{z}_{\rm a}(T)$ and $\overline{S}^{z}_{\rm d}(T)$. As a result, within the framework of MFA, YIG can be regarded as a two-sublattice ferrimagnet.
@@
Here ${\cal J}_{\rm aa}\equiv J_{\rm aa}Z_{\rm aa}$, ${\cal J}_{\rm dd}\equiv J_{\rm dd}Z_{\rm dd}$, and ${\cal J}_{\rm ad}\equiv J_{\rm ad}Z_{\rm ad}$, ${\cal J}_{\rm da}\equiv J_{\rm ad}Z_{\rm da}$. Here, we recall that $Z_{ij}$ are the coordination numbers $Z_{\rm ad}=6$, $Z_{\rm da}=4$, $Z_{\rm aa}=8$, and $Z_{\rm dd}=4$. When combined with the experimental values of the exchange integrals (7), the resulting values are:
@@
Equations (13) lead to a natural generalization of the mean field equation (4c):
@@
$\overline{S}_{\rm d}^{z}(T)=\,\,$ $b_{{}_{S}}\Big{[}\frac{\overline{S}_{\rm a}^{z}(T){\cal J}_{\rm ad}+\overline{S}_{\rm d}^{z}(T){\cal J}_{\rm dd}}{T}\Big{]}\ .$ (15b)
Employing the expansion (5), we solve $\overline{S}_{\rm a}^{z}=\overline{S}_{\rm d}^{z}=0$, to find the Curie temperature $T_{{}_{\rm C}}^{{}^{\rm MF}}$ within the MFA:
@@
One can show that Eqs. (15) results in the square-root scaling of $\overline{S}_{\rm d}^{z}$ and $\overline{S}_{\rm a}^{z}$ as $T$ approaches $T_{{}_{\rm C}}^{{}^{\rm MF}}$ from below:
$\overline{S}_{\rm a}^{z}\propto\sqrt{T_{{}_{\rm C}}^{{}^{\rm MF}}-T}\,,\quad\overline{S}_{\rm a}^{z}\propto\sqrt{T_{{}_{\rm C}}^{{}^{\rm MF}}-T}\,,$ (18)
@@
Consider now the solution of Eqs. (15) when the temperature approaches zero from above $T\to 0^{+}$. In this limit, the argument of $b_{{}_{S}}$ in (15a) approaches $\infty$- ∞ (assuming that $\overline{S}_{\rm d}^{z}>0$, $J_{\rm ad}<0$, and $|J_{\rm ad}|\gg|J_{\rm aa}|$) and approaches +$\infty$+ ∞ in (15b) (if $\overline{S}_{\rm a}^{z}<0$ and $|J_{\rm ad}|\gg|J_{\rm dd}|$). Using the equation for $b_{{}_{S}}(x)$ in the limits $|x|\gg 1$
$\begin{split}b_{{}_{S}}(x)&\approx S-\exp(-x)\,,\quad x\to+\infty\,,\\ b_{{}_{S}}(x)&\approx-S+\exp(x)\,,\quad x\to-\infty\,,\end{split}$ (19a)
@@
We see that as $T\to 0$, $\overline{S}_{\rm a}^{z}$ approaches $-S$, while $\overline{S}_{\rm d}^{z}$ approaches $+S$. At small $T$, the deviation of both $|\overline{S}_{\rm a}^{z}|$ and $|\overline{S}_{\rm a}^{z}|$ from $S$ is exponentially small.
@@
Green dashed and solid lines ⑨ and ⑩ represent the normalized spins of sublattice (a) $-\overline{S}^{z}_{\mathrm{a}}(T)/\overline{S}^{z}_{\mathrm{a}}(0)$, and sublattice (d) $\overline{S}^{z}_{\mathrm{d}}(T)/\overline{S}^{z}_{\mathrm{d}}(0)$ respectively, defined by Eqs. (15) based on exchange integrals from [23]. Vertical dashed lines mark the Curier temperatures $T_{{}_{\rm C}}^{\rm Exp}$ (black) and $T_{{}_{\rm C}}^{\rm MF}$ (red).
@@
Fig. 4 presents the results of the numerical solutions of MF Eqs. (15). Solid blue line ⑤, also shown in Fig. 2, represents the normalized total spin in YIG, defined by Eq. (8) with $N_{\rm d}=12$ and $N_{\rm a}=8$. The green dashed and solid lines ⑨ and ⑩ correspond to the normalized spins of the sublattice (a) $-\overline{S}^{z}_{\mathrm{a}}(T)/\overline{S}^{z}_{\mathrm{a}}(0)$, and sublattice (d) $\overline{S}^{z}_{\mathrm{d}}(T)/\overline{S}^{z}_{\mathrm{d}}(0)$, respectively. Here, in (15) we used the exchange integrals (14).

### v2

# Unified theory of magnetization temperature dependence in ferrimagnets

@@ Abstract
Recent advancements in spintronics and fundamental physical research have brought increased attention to the rare-earth-based magnetically ordered materials. One of the important properties of these materials is the temperature dependence of the spontaneous magnetization $M(T)$. Recently, a successful framework was proposed for the theoretical description of $M(T)$ across the entire temperature range from $T=0$ K to the Curie temperature $T_{{}_{\rm C}}$ in simple cubic ferromagnetics, EuO and EuS. We extend this approach to compute and analyze $M(T)$ for a more complex magnetic material: multi-sublattice collinear ferrimagnets such as Yttrium Iron Garnet, (YIG) Y3Fe5O12. YIG is a unique material that has a high Curie temperature well above room temperature. It features a cubic crystallographic structure characterized by low anisotropy and minimal spin-orbit interaction. These properties result in an exceptionally low spin-wave damping, which is important for fundamental studies as well as engineering applications. The primary challenge in describing ferrimagnets theoretically arises from its intricate magnetic structure, characterized by multiple magnetic sublattices. We analyzed and generalized for multi-sublattice collinear ferrimagnets two well-known approximations describing $M(T)$. The first approach is the Bloch $\frac{3}{2}$-law, which describes the suppression of $M(T)$ due to spin-wave excitation, and is valid in the low-temperature limit $T\ll T_{{}_{\rm C}}$. The second one is Weiss’s mean-field approximation, which provides a reasonable description of $M(T)$ near $T_{{}_{\rm C}}$. Using a single tuning parameter, we combine these two approaches to describe $M(T)$ for any $0\leq T\leq T_{{}_{\rm C}}$. The theoretical result for $M(T)$ aligns well with our measurements and the previously available experimental data across the entire temperature range. We also demonstrate that experimental and theoretical dependencies $M(T)$ follow the mean-field prediction $\sqrt{T_{{}_{\rm C}}-T}$ for almost all temperatures.
[…]
@@ I Introduction
Promising perspectives for engineering applications revived the attention to fundamental studies of the physics of magnetic materials. To our understanding, the first step in developing a consistent theory of magnetic dielectrics is the study of their most fundamental properties: the temperature dependence of spontaneous magnetization $M(T)$ in thermodynamic equilibrium. The magnetization $M(T)$ plays a central role in the theories concerning the temperature dependence of other aspects of magnetodielectrics, such as magnetic susceptibilities, heat capacities, spin-wave frequency spectra, matrix elements of the magnetic dipole-dipole and exchange interactions, damping frequencies of magnons, and even quantum corrections to $M(T)$.
Recently [6], the behavior of $M(T)$ was studied for the simplest possible magnetic system: cubic ferromagnetic crystals containing only one magnetic ion per elementary cell. Currently, the only known and experimentally studied example of such a system is body-centered crystals of EuO and EuS.
The analysis in [6] is based on the well-known Weiss mean-field approximation (MFA) [7] with the Heisenberg exchange interaction [8]. In the MFA, an effective magnetic field acts on each magnetic ion due to exchange interactions with its nearest neighbors. The larger their coordination number $Z$ (i.e., the number of nearest neighbors), the smaller the fluctuations of this field, and the better the MFA performs. In the limit $Z\to\infty$, MFA becomes exact. In EuO, $Z=12\gg 1$ and the resulting value of the Curie temperature for EuO $T^{{}^{\rm MF}}_{{}_{\rm C}}\approx 86.6\,$K, predicted by MFA, is only about 20% larger than its experimental value $T^{\rm Expt}_{{}_{\rm C}}\approx 69.8\,$K.
@@
A natural further step in the development of the theory of the spontaneous magnetization is to extend it to multi-sublattice ferri- and antiferromagnets. This is the objective of the present paper. To make our study more concrete and provide a basis for comparing our theory with experiments, we have chosen a specific example: the twenty-sublattice ferrimagnetic Yttrium Iron Garnet (YIG), Y3Fe5O12. YIG is one of the most important materials in microwave magnetics. In physics of magnetically ordered crystals, it plays a crucial role much like the role of Drosophila melanogaster (fruit fly) in genetics [9] or water in hydrodynamics  [10].
@@
Various experimental datasets and theoretical results are labeled as follows. Experimental data: blue dots ① – 1964-Anderson results [20], orange dots ② – 1974 data Hansen and others [21], black solid line ③ – our current results of a 1 mm YIG sphere magnetized in the ⟨111⟩ direction were measured via VSM.
Violet dashed line ④ – all-temperature fit of experimental data ② $\overline{S}^{z}=A(T_{{}_{\rm C}}-T)^{\beta}$ with $A$ = 1.045 and $\beta\simeq 0.50\pm 0.005$; for details see Appendix H.
@@
Our aim in this paper is to improve the description of multi-sublattice ferrimagnets, focusing on a theory of spontaneous magnetization $M(T)$ over an entire temperature range from $T=0$ to the Curie temperature $T_{{}_{\rm C}}$.
@@
We begin our study by presenting in Fig. 2 the earlier available experimental data [20, 21] for $\overline{S}^{z}(T)/\overline{S}^{z}(0)$, as blue and orange dots labeled as ① and ② and our own accurate measurements down to temperatures near 2 K, marked as the black solid line ③. The experimental Curie temperature $T_{{}_{\rm C}}^{\rm Expt}\approx 560\,$K is clearly visible in the figure. A small tail above $T_{{}_{\rm C}}^{\rm Expt}$ originates from the external magnetic field (about 10 kOe) used in experiments to suppress the domain structure of YIG, interfering with measurements.
@@
We fit the experimental data with the $\overline{S}^{z}(T)\propto(T_{{}_{\text{C}}}-T)^{\beta}$ with $\beta=0.50\pm 0.005$, shown in Fig. 2 as a violet dashed line ④. A detailed description of the fitting method can be found in Appendix H. Unexpectedly, the normal scaling agrees well with experiments conducted not only near the critical temperature $T_{{}_{\text{C}}}$ but across nearly all temperatures ranging from approximately $T_{{}_{\rm C}}^{\rm Expt}\approx 560\,\text{K}$ down to about $100\,\text{K}$, where $\overline{S}^{z}(T)/\overline{S}^{z}(0)\approx 0.95$. This behavior is in sharp contrast with the anomalous scaling of $M(T)$ observed in simple ferromagnets like EuO and EuS, where the exponent $\beta$ is approximately $1/3$ [6].
@@
To this end, in Sec. III.1, we consider the MFA for two-sublattice ferrimagnets and derive Eq. (15) that describe it. Next, we derived an analytical expression for the Curie temperature, Eq. (16a), along with the ratio of the sublattice magnetizations near $T_{{}_{\rm C}}$, Eq. (16b). The numerical solution of Eq. (15) in the form Eq. (8) is plotted in Fig. 2 by the blue solid line ⑤. As expected, near $T_{{}_{\rm C}}^{{}^{\rm MF}}$ it demonstrates normal scaling behavior $\overline{S}^{z}(T)\propto\sqrt{T_{{}_{\rm C}}^{{}^{\rm MF}}-T}$, denoted by the blue dashed line ⑥.
The resulting Curie temperature $T_{{}_{\rm C}}^{{}^{\rm MF}}\approx 894\,$K calculated using values of the YIG exchange integrals [23] exceeds its experimental value $T_{{}_{\rm C}}^{\rm Expt}=560\,$K almost twice, while in EuO the discrepancy is only about 20%. This is discouraging but not unexpected, since the dominant inter-sublattice interactions in YIG are carried through only $Z_{\rm YIG}\approx 5$ neighbors on average. Therefore, the applicability parameter $1/Z_{\rm YIG}\approx 1/5$ is larger than $1/Z_{\rm EuO}\lesssim 1/12$, and poorer performance of MFA compared to EuO is expected.
@@
This issue is addressed in Sec. III.2 where we formulated spin-wave approximation (SWA) (21) for suppression of magnetization in multi-sublattice magnetics. To calculate the magnitization in the framework of SWA, an explicit knowledge of the magnon frequency spectra is required.
It is well known that the frequency spectra of the magnetic material with $n$ sublattices involve $n$ branches. For collinear magnetic structures with $n_{+}$ sublattices oriented along the external magnetic field and $n_{-}=n-n_{+}\leq n_{+}$ sublattices oriented in the opposite direction, the total $n$ branches can be categorized into two groups: $n_{+}$ “ferromagnetic” modes and $n_{-}$ “antiferromagnetic” modes. According to Eq. (21), the excitation of a single ferromagnetic magnon reduces the total $z$-projection of spin, $\overline{S}^{z}$, by unity. In contrast, the excitation of a single antiferromagnetic magnon increases $\overline{S}^{z}$ by unity.
@@
Formally speaking, Eq. (21) is exact in the low-temperature range, say for $T\ll T_{{}_{\rm C}}$. To our surprise, the numerical solution of Eq. (21), represented by the green solid line ⑦ in Fig. 2, aligns well with the experimental data ①, ②, and ③ from $T=0$ up to $T\approx 400\,$K. At higher temperatures $\overline{S}^{z}(T)$ decreases almost linearly, crossing zero at $T_{\times}\approx 770\,$K and then formally becomes negative.
In Eq. (21), we assumed temperature-independent frequency dispersion relations taken at T=0T=0\,K. However, the frequency spectra depend on $T$ according to Eq. (25), where a number of additional assumptions were made. When this temperature dependence is taken into account in Eq. (21), the resulting numerical solution, which is still correct at low $T$, is very strongly suppressed with the increasing temperature, crossing now zero at $T_{\times}\approx 410\,$K.
@@
In the simple case of a one-sublattice ferromagnet, this concept was demonstrated in Ref. [6] and briefly reminded in Sec. IV.1 for completeness. Its generalization for multi-sublattice ferrimagnets like YIG is not straightforward and requires significant efforts, including the application of the diagrammatic perturbation approach for spin operators, such as the Belinicher-Lvov diagrammatic technique [24]. This approach is beyond the scope of the current paper.
@@
The numerical solution of Eqs. (27) and (30) with $\delta=0.428$ is shown in Fig. 2 as the red solid line ⑧. Bearing in mind that we computed the solution ⑧ using the single tuning parameter $\delta$, we consider its agreement with experimental data for $\overline{S}^{z}(T)$ across the entire temperature range as fully satisfactory. It is worth noting that the theoretical dependence ⑧ of $\overline{S}^{z}$ retains normal scaling over $100\lesssim T\lesssim T_{\rm{}_{C}}$, much wider than the MFA.
As previously discussed, the method for studying the temperature dependence of magnetization in single-sublattice ferromagnetics like EuO and EuS, as presented in [6], shares several similarities with the approach used for multi-sublattice ferrimagnets, such as YIG, developed in this paper. While the initial steps are similar, they diverge significantly in the final stages. Details of their comparison are provided in Appendix B.
We summarize our findings in Sec.V.
@@ II.1 Weiss-Heisenberg mean-field approximation
The analysis of magnetic properties in magnetodielectrics begins with the Weiss mean-field approximation [7] (MFA) with the Heisenberg exchange Hamiltonian [8]. In our paper, we adopt the simplest form of this Hamiltonian, which describes the interaction between the spins $\bm{S}_{i}$and $\bm{S}_{j}$, located at positions $\bm{r}_{i}$ and $\bm{r}_{j}$ respectively:
@@
Here, exchange integrals $J_{ij}$ represent the energy associated with the exchange interactions. In addition to exchange interactions, much weaker magnetic dipole-dipole interactions may also play some role in the very vicinity of the Curie temperature  [25, 26, 27]. The dipole-dipole interactions determine the frequency spectrum of long spin-waves. We discuss their contribution in Appendices E and F.
@@
$~\thinspace b_{{}_{S}}(x)$ $=$ $\Big(S+\frac{1}{2}\Big)\coth\!\Big[\Big(S+\frac{1}{2}\Big)x\Big]-\frac{1}{2}\coth\Big(\frac{x}{2}\Big)\ .$ (4b)
@@ II.2 Crystallographic and magnetic structure of YIG
[…]
@@
Over several decades, a variety of exchange integrals for YIG have been proposed, some of which are listed in Tab. 1 and discussed in Appendix D. For this paper, we adopt the following values:
@@ II.3 Spin-waves spectra of YIG
[…]
Recall that in YIG $n_{+}=12$, $n_{-}=8$, and $n=n_{+}+n_{-}=20$. The explicit and quite cumbersome equations for $A_{ij}(\bm{k})$, $B_{ij}(\bm{k})$ and $D_{ij}(\bm{k})$ one finds in [23].
@@ III.1 MFA for two-sublattice ferrimagnets
As discussed in the Introduction, the MFA focuses solely on the mean values of the spin $\overline{S}^{z}_{j}$ at site $\bm{r}_{j}$. In YIG, there are only two crystallographically distinct positions of Fe ions, labeled (a) and (d), with mean spins $\overline{S}^{z}_{\rm a}(T)$ and $\overline{S}^{z}_{\rm d}(T)$. As a result, within the framework of MFA, YIG can be regarded as a two-sublattice ferrimagnet.
@@
Here ${\cal J}_{\rm aa}\equiv J_{\rm aa}Z_{\rm aa}$, ${\cal J}_{\rm dd}\equiv J_{\rm dd}Z_{\rm dd}$, and ${\cal J}_{\rm ad}\equiv J_{\rm ad}Z_{\rm ad}$, ${\cal J}_{\rm da}\equiv J_{\rm ad}Z_{\rm da}$. Here, we recall that $Z_{ij}$ are the coordination numbers $Z_{\rm ad}=6$, $Z_{\rm da}=4$, $Z_{\rm aa}=8$, and $Z_{\rm dd}=4$. When combined with the experimental values of the exchange integrals (7), the resulting values are
@@
Equations (13) lead to a natural generalization of the mean field equation (4c)
@@
$\overline{S}_{\rm d}^{z}(T)=\,\,$ $b_{{}_{S}}\Big[\frac{\overline{S}_{\rm a}^{z}(T){\cal J}_{\rm da}+\overline{S}_{\rm d}^{z}(T){\cal J}_{\rm dd}}{T}\Big]\ .$ (15b)
Employing the expansion (5), we solve $\overline{S}_{\rm a}^{z}=\overline{S}_{\rm d}^{z}=0$, to find the Curie temperature $T_{{}_{\rm C}}^{{}^{\rm MF}}$ within the MFA
@@
One can show that Eqs. (15) results in the square-root scaling of $\overline{S}_{\rm d}^{z}$ and $\overline{S}_{\rm a}^{z}$ as $T$ approaches $T_{{}_{\rm C}}^{{}^{\rm MF}}$ from below
$\overline{S}_{\rm a}^{z}\propto\sqrt{T_{{}_{\rm C}}^{{}^{\rm MF}}-T}\,,\quad\overline{S}_{\rm d}^{z}\propto\sqrt{T_{{}_{\rm C}}^{{}^{\rm MF}}-T}\,$, (18)
@@
Consider now the solution of Eqs. (15) when the temperature approaches zero from above $T\to 0^{+}$. In this limit, the argument of $b_{{}_{S}}$ in (15a) approaches -$\infty$ (assuming that $\overline{S}_{\rm d}^{z}>0$, $J_{\rm ad}<0$, and $|J_{\rm ad}|\gg|J_{\rm aa}|$) and approaches $\infty$ in (15b) (if $\overline{S}_{\rm a}^{z}<0$ and $|J_{\rm ad}|\gg|J_{\rm dd}|$). Using the equation for $b_{{}_{S}}(x)$ in the limits $|x|\gg 1$,
[…]
@@
We see that as $T\to 0$, $\overline{S}_{\rm a}^{z}$ approaches S-S, while $\overline{S}_{\rm d}^{z}$ approaches +S+S. At small $T$, the deviation of both $|\overline{S}_{\rm a}^{z}|$ and $|\overline{S}_{\rm a}^{z}|$ from $S$ is exponentially small.
@@
Green dashed and solid lines ⑨ and ⑩ represent the normalized spins of sublattice (a) $\overline{S}^{z}_{\mathrm{a}}(T)/\overline{S}^{z}_{\mathrm{a}}(0)$, and sublattice (d) $\overline{S}^{z}_{\mathrm{d}}(T)/\overline{S}^{z}_{\mathrm{d}}(0)$ respectively, defined by Eqs. (15) based on exchange integrals from [23]. Vertical dashed lines mark the Curier temperatures $T_{{}_{\rm C}}^{\rm Exp}$ (black) and $T_{{}_{\rm C}}^{\rm MF}$ (red).
@@
Fig. 4 presents the results of the numerical solutions of MF Eqs. (15). Solid blue line ⑤, also shown in Fig. 2, represents the normalized total spin in YIG, defined by Eq. (8) with $N_{\rm d}=12$ and $N_{\rm a}=8$. The green dashed and solid lines ⑨ and ⑩ correspond to the normalized spins of the sublattice (a) $\overline{S}^{z}_{\mathrm{a}}(T)/\overline{S}^{z}_{\mathrm{a}}(0)$, and sublattice (d) $\overline{S}^{z}_{\mathrm{d}}(T)/\overline{S}^{z}_{\mathrm{d}}(0)$, respectively. Here, in (15) we used the exchange integrals (14).